# Espresso: Robust Concept Filtering in Text-to-Image Models

### v2

@@ Abstract.
Diffusion based text-to-image (T2I) models generate high fidelity images for given textual prompts. They are trained on large datasets scraped from the Internet, potentially containing unacceptable concepts (e.g., copyright infringing or unsafe). Retraining T2I models after filtering out unacceptable concepts in the training data is inefficient and degrades utility. Hence, there is a need for concept removal techniques (CRTs) which are i) effectivein removing unacceptable concepts ii) utility-preservingon acceptable concepts, and iii) robustagainst evasion with adversarial prompts. None of the prior CRTs satisfy all these requirements simultaneously.
@@
We introduce Espresso, the first robust concept filter based on Contrastive Language-Image Pre-Training (CLIP). It identifies unacceptable concepts by projecting the generated image’s embedding onto the vector connecting both unacceptable and acceptable concepts in the joint text-image embedding space. This ensures robustness by restricting the adversary to adding noise only along this vector, in the direction of the acceptable concept. Further fine-tuning Espresso to separate embeddings of acceptable and unacceptable concepts, while preserving their pairing with image embeddings, ensures both effectiveness and utility. We evaluate Espresso on eleven concepts to show that it is effective ($\sim$5CLIP accuracy on unacceptable concepts), utility-preserving ($\sim$93normalized CLIP score on acceptable concepts), and robust ($\sim$4CLIP accuracy on adversarial prompts for unacceptable concepts). Finally, we present theoretical bounds for certified robustness of Espresso against adversarial prompts, and an empirical analysis.
copyright: none
1. Introduction
@@
Diffusion based text-to-image (T2I) models have demonstrated a remarkable ability to generate high quality images from textual prompts (Rombach et al., 2022a, b; Ramesh et al., 2021). These models are trained on large datasets of unfiltered content from the Internet (Radford et al., 2021; Schuhmann et al., 2022). Due to their large capacity, T2I models memorize specific concepts, as seen in the generated images (Korn, 2023; Somepalli et al., 2023; Carlini et al., 2023a). Some of these concepts, may be unacceptable for various reasons such as copyright infringement (e.g., a movie character or celebrity) or inappropriateness (e.g., “nudity” or “violence”) (Gandikota et al., 2023a; Schramowski et al., 2023; Heng and Soh, 2023b). Retraining T2I models after filtering out unacceptable concepts is inefficient, only partially effective (Li et al., 2024; Rombach et al., 2023), and compromises utility (Gandikota et al., 2023a). Hence, there is a need for concept removal techniques (CRTs) to minimize unacceptable concepts in generated images.
@@
Ideally, CRTs should be effective in reducing the generation of unacceptable concepts while preserving the utility on all others, and robust to evasion with adversarial prompts. As we show in Section 6, none of the existing CRTs simultaneously satisfy these requirements: i) filtering CRTs, which detect unacceptable concepts, lack robustness ((Rando et al., 2022) and Section 6.2) ii) fine-tuning CRTswhich modify T2I models, trade-off effectiveness and utility (Gandikota et al., 2023a; Kumari et al., 2023a; Heng and Soh, 2023b; Schramowski et al., 2023), and lack robustness (Tsai et al., 2024; Wen et al., 2023; Pham et al., 2024a; Yang et al., 2024b). Designing a CRT meeting all the requirements is an open problem. Our goal is to design such a CRT.
@@
We opt to use a filter as it will not alter the T2I model, thus reducing its impact on utility. We construct our filter using the Contrastive Language-Image Pre-Training (CLIP) (Radford et al., 2021), an essential component of T2I models. CLIP co-locates the embeddings of textual prompts and their corresponding images within a unified space. CLIP is pre-trained on a vast dataset which encodes a broad spectrum of concepts (Radford et al., 2021), making it a versatile choice for a filter, unlike specialized classifiers (e.g., (Qu et al., 2023b)). However, relying on CLIP to identify unacceptable images by solely measuring the distance between the embeddings of a generated image, and an unacceptable concept, was shown not to be robust ( (Rando et al., 2022)). Subsequently, several prior works have identified filtering as a viable direction which needs further exploration (Rando et al., 2022; Li et al., 2024).
@@
We present a robust content filter, Espresso, that identifies unacceptable concepts in generated images using the distance of their embeddings to the text embeddings of both acceptable and unacceptable concepts. We project the image embedding onto the vector connecting the text embeddings of acceptable and unacceptable concepts, which makes it harder to generate effective adversarial prompts: the adversary is restricted to adding noise only along this vector, in the direction of the acceptable concept. Fine-tuning Espresso further increases the separation between the text embeddings of acceptable and unacceptable concepts for effectiveness, while maintaining their pairing with their corresponding image embeddings for utility. This fine-tuning approach is further supported by fine-tuning CRTs which show that such an optimization leads to better effectiveness (Kumari et al., 2023a; Tsai et al., 2024; Gal et al., 2022). We claim the following main contributions: we present
  1. (1)
@@
    Espresso  [^1] We will open-source the code upon publication., the first robust (CLIP-based) content filter which identifies unacceptable concepts by measuring the distance of generated images’ embeddings to the text embeddings of both unacceptable and acceptable concepts (Section 4),
  2. (2)
@@
    a comprehensive comparative evaluation (Section 5) of the fine-tuned variant of Espresso with six state-of-the-art fine-tuning based CRTs, and one filtering CRT, showing that Espresso is effective ($\sim$5CLIP accuracy on unacceptable concepts), utility-preserving ($\sim$93normalized CLIP score on acceptable concepts), and robust ($\sim$4CLIP accuracy of on adversarial prompts) (Section 6), and
  3. (3)
@@
2. Background
We describe T2I models (Section 2.1), different concept removal mechanisms (Section 2.2), and attacks against them (Section 2.3).
2.1. Diffusion based T2I Models
@@
where $\tau$ is the temperature parameter for scaling the predictions (Radford et al., 2021; Liang et al., 2022).
@@
Several prominent T2I models (e.g., Stable Diffusion v1.4 (SDv1.4)) improve the efficiency of the diffusion process by conducting it in the embedding space of a variational autoencoder (VAE), rather than on images (Rombach et al., 2022a). For a VAE decoder $\mathcal{D}$, VAE encoder $\mathcal{E}$, and $z_{t}\in\mathcal{E}(\texttt{$x$})$ as the latent representation of $x$ in the VAE’s latent space, $\epsilon_{\theta}$ is trained by minimizing the following objective:
$\mathcal{L}=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,1),z_{t}\in\mathcal{E}(\texttt{$x$}),\phi_{p}(\texttt{p}),t}[||\epsilon-\epsilon_{\theta}(z_{t},\phi_{p}(\texttt{p}),t)||_{2}^{2}]$. The final image is generated from the approximation, $\tilde{z_{0}}$, by passing it through $\mathcal{D}$: $\tilde{x_{0}}=\mathcal{D}(\tilde{z_{0}})$. Table 1 summarizes frequently used notations.
Table 1. Frequently used notations and their descriptions.
Notation Description
T2I Text-to-Image
$\phi$ Text encoder (e.g., CLIP)
CRT Concept Removal Technique
$\mathcal{A}dv$ Adversary
$x$ Generated image
$x^{u}$ Generated image with unacceptable concept
$x^{a}$ Generated image with acceptable concept
$\phi_{x}(x)$ CLIP embedding for $x$
@@
p Textual prompt
$c^{a}$ Phrase for acceptable concept
$c^{u}$ Phrase for unacceptable concept
$\texttt{p}^{u}$ Textual prompt containing $c^{u}$
$\texttt{p}^{a}$ Textual prompt containing $c^{a}$
$\phi_{p}(\texttt{p})$ CLIP embedding for textual prompt p
$\texttt{p}^{adv}$ Adversarially generated textual prompt
$\mathcal{D}_{te}^{a}$ Test dataset with acceptable prompts/images
$\mathcal{D}_{te}^{u}$ Test dataset with unacceptable prompts/images
$\mathcal{D}_{val}^{a}$ Validation dataset with acceptable prompts/images
$\mathcal{D}_{val}^{u}$ Validation dataset w/ unacceptable prompts/images
$\mathcal{D}_{adv}^{u}$ Test dataset with adversarial prompts/images
f T2I model with CRT where f: $c\rightarrow x$
F Function for Espresso classifier
$\alpha$ Regularization parameter for Espresso fine-tuning
$\overline{\cdot}$ Normalization function
$\epsilon_{\theta}$ Diffusion model parameterized by $\theta$
$\tau$ Temperature parameter
2.2. Concept Removal Techniques
The textual phrase for an acceptable concept is $c^{a}$, and for an unacceptable concept is $c^{u}$. An image $x$ generated from a T2I model may either contain an unacceptable concept (referred to as $x^{u}$) or an acceptable one (referred to as $x^{a}$). Similarly, a text prompt p may contain a phrase for an acceptable concept ($\texttt{p}^{a}$) or an unacceptable concept ($\texttt{p}^{u}$). An example of an unacceptable prompt $\texttt{p}^{u}$ containing an unacceptable concept $c^{u}$, Captain Marvel is “Captain Marvel soaring through the sky”. CRTs seek to thwart the generation of $x^{u}$ by either fine-tuning the T2I model to suppress $x^{u}$, or using a classifier as a filter to detect $x^{u}$. We first present six state-of-the-art fine-tuning CRTs:
Concept Ablation (CA(Kumari et al., 2023a) fine-tunes the T2I model to minimize the KL divergence between the model’s output for $\texttt{p}^{u}$ and $\texttt{p}^{a}$ to force the generation of $x^{a}$ instead of $x^{u}$. Formally, they optimize the following objective function:
@@ 2.2. Concept Removal Techniques
$\begin{split}\mathcal{L}_{\texttt{CA}}=\mathbb{E}_{\epsilon,z_{t},c^{u},c^{a},t}[w_{t}||\epsilon_{\theta}(z_{t},\phi_{p}(\texttt{p}^{a}),t).sg()-\epsilon_{\theta}(z_{t},\phi_{p}(\texttt{p}^{u}),t)||_{2}^{2}]\end{split}$
@@
Forget-Me-Not (FMN(Zhang et al., 2023b) minimizes the activation maps for $c^{u}$ by modifying $\epsilon_{\theta}$’s cross-attention layers. Further, fine-tuning $\epsilon_{\theta}$, instead of just the cross-attention layers, results in degraded utility.
Selective Amnesia (SA(Heng and Soh, 2023b) fine-tunes T2I models by adapting continuous learning techniques (elastic weight consolidation and generative replay) for T2I models to forget a concept. They optimize $\mathbb{P}(\texttt{$x$}|\theta^{*},c^{u})$, the probability of generating $x$ given $\theta^{*}$ and $c^{u}$, where $\theta^{*}$ are the frozen parameters of the original T2I model:
$\begin{split}\mathcal{L}_{\texttt{SA}}=-\mathbb{E}_{\mathbb{P}(\texttt{$x$}|\texttt{p})\mathbb{P}_{f}(c^{u})}[\text{log}\mathbb{P}(\texttt{$x$}|\theta^{*},c^{u})]-\lambda\sum_{i}\frac{M_{i}}{2}(\theta_{i}^{*}-\theta_{i})^{2}+\\ \mathbb{E}_{\mathbb{P}(\texttt{$x$}|\texttt{p})\mathbb{P}_{r}(c^{a})}[\text{log}\mathbb{P}(\texttt{$x$}|\theta,c^{a})]\end{split}$
Erased Stable Diffusion (ESD(Gandikota et al., 2023a) fine-tunes the T2I model by modifying the reverse diffusion process to reduce the probability of generating $x^{u}$:
$\epsilon_{\theta}(\texttt{$x$}_{t},\phi_{p}(c^{u}),t)=\epsilon_{\theta^{*}}(\texttt{$x$}_{t},t)+\eta(\epsilon_{\theta^{*}}(\texttt{$x$}_{t},\phi_{p}(c^{u}),t)-\epsilon_{\theta^{*}}(\texttt{$x$}_{t},t))$
Unified Concept Editing (UCE(Gandikota et al., 2024) fine-tunes the T2I model’s cross-attention layers with the help of a language model to minimize the influence of $c^{u}$, while keeping the influence of remaining concepts unchanged. Their fine-tuning objective is
$\begin{split}\mathcal{L}_{\texttt{UCE}}=\sum_{c^{u}\in\mathcal{C}^{u},c^{a}\in\mathcal{C}^{a}}||Wc^{u}-W^{\text{old}}c^{a}||_{2}^{2}+\sum_{c\in S}||Wc-W^{\text{old}}c||_{2}^{2}\end{split}$
where $W$, $W^{\text{old}}$ are the parameters of the fine-tuned and original cross-attention layers in $\epsilon_{\theta}$, $\mathcal{C}^{u}$ and $\mathcal{C}^{a}$ are the space over some pre-defined unacceptable and acceptable concepts, and $S$ is a set of concepts for which to preserve utility.
Safe diffusion (SDD(Kim et al., 2023) fine-tunes the T2I model by encouraging the diffusion model noise conditioned on $c^{u}$ to match the unconditioned noise, while minimizing the utility drop using the following objective function:
$\begin{split}\mathcal{L}_{\texttt{SDD}}=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,1),z_{t}\in\mathcal{E},c^{u},t}[||\epsilon_{\theta}(z_{t},\phi_{p}(c^{u}),t)-\epsilon_{\theta}(z_{t},t).sg()||_{2}^{2}]\end{split}$
We now describe two filtering CRTs:
Stable Diffusion Filter (SD-Filter) (Radford et al., 2021) is black-box and the design of the filter is not publicly available. However, Rando et. al. (Rando et al., 2022) hypothesize that it involves computing the cosine similarity between the embeddings of a generated image, $x$, and a pre-defined set of $c^{u}$. If the cosine similarity is greater than some threshold ($\Gamma$), then the unacceptable concept is detected. Formally, their filter $\texttt{F}_{SD}$ can be described as
where $\text{cos}(\phi_{x}(\texttt{$x$}),\phi_{p}(c^{u}))=\overline{\phi_{x}(\texttt{$x$})}\cdot\overline{\phi_{p}(c^{u})}$, and $\overline{\cdot}$ denotes normalization. Here $\texttt{F}_{SD}(\texttt{$x$})=0$ indicates $x^{a}$ and $\texttt{F}_{SD}(\texttt{$x$})=0$ indicates $x^{u}$. Note that $\Gamma$ is possibly different for different $c^{u}$.
Unsafe Diffusion (UD(Qu et al., 2023b) is the current state-of-the-art filtering CRT which trains a multi-headed neural network classifier on top of CLIP to identify $x^{u}$. The different heads for the neural network are for classifying different $c^{u}$: nudity, violence, disturbing, hateful, and political. Their objective is given as $\texttt{F}_{\texttt{UD}}(\texttt{$x$})=\text{MLP}(\phi_{x}(\texttt{$x$}))$ where MLP is a multi-layer perceptron, and $\texttt{F}_{\texttt{UD}}\in\{0,1\}$.
2.3. Evading Concept Removal Techniques
An adversary ($\mathcal{A}dv$) may generate adversarial prompts ($\texttt{p}^{adv}$) to evade CRTs and force the T2I model to generate unacceptable images. We denote a dataset containing adversarial prompts and their corresponding images as $\mathcal{D}_{adv}^{u}=\{(\texttt{$x$}^{adv}_{j},\texttt{p}^{adv}_{j})\}_{j=1}^{N}$. $\mathcal{A}dv$ is assumed to account for the target T2I model, f, using a CRT. A dumb $\mathcal{A}dv$ who does not account for this is naïve. All these attacks generate $\texttt{p}^{adv}$ as an optimization problem using some reference $x^{u}$ (or the difference between $\texttt{p}^{a}$ and $\texttt{p}^{u}$) as the ground truth. Different attacks vary in ways to solve this optimization. We present four state-of-the-art attacks below:
PEZ (Wen et al., 2023) generates $\texttt{p}^{adv}$ by identifying text tokens which minimize the following objective function: $\mathcal{L}_{PEZ}=1-\text{cos}(\phi_{p}(\texttt{p}^{adv}),\phi_{x}(\texttt{$x^{u}$}))$.
RingBell (Tsai et al., 2024) uses a genetic algorithm to identify tokens for $\texttt{p}^{adv}$ by minimizing the difference between the embedding vectors of $\texttt{p}^{u}$ and $\texttt{p}^{a}$: $\phi_{p}(\hat{\texttt{p}})=\frac{1}{N}\sum_{i=1}^{N}\{\phi_{p}(\texttt{p}^{u}_{i})-\phi_{p}(\texttt{p}^{a}_{i})\}$ from a set of $N$ prompts. The optimization can be formulated as: $\text{min}_{\texttt{p}^{adv}}||\phi_{p}(\texttt{p}^{adv})-(\phi_{p}(\texttt{p}^{init})+\eta\cdot\phi_{p}(\hat{\texttt{p}}))||^{2}$ where $\texttt{p}^{init}$ is an initial target prompt which is modified into $\texttt{p}^{adv}$. Further, the optimization relies on a fitness function using mean-squared error: $\mathcal{L}_{RingBell}=\mathcal{L}_{MSE}(\phi_{p}(c^{u}),\phi_{p}(\texttt{p}^{adv}))$ where $c^{u}$ is a prompt that generates $x^{u}$ in a T2I model.
@@ 2.3. Evading Concept Removal Techniques
CCE (Pham et al., 2024a) uses textual-inversion to generate $\texttt{p}^{adv}$ (Gal et al., 2023). $\mathcal{A}dv$ updates CLIP’s vocabulary to include a token new “¡s¿” which, when included in $\texttt{p}^{adv}$, generates $\mathcal{A}dv$’s desired image. To optimize ¡s¿, we find $v$, an embedding for ¡s¿, corresponding to least loss $\mathcal{L}_{CCE}=\mathcal{L}_{MSE}(\epsilon,\epsilon_{\theta}(z_{t},v,t))$, where $\epsilon\sim\mathcal{N}(0,1)$.

### v4

@@ Abstract
Diffusion based text-to-image (T2I) models generate high fidelity images for given textual prompts. They are trained on large datasets scraped from the Internet, potentially containing unacceptable concepts (e.g., copyright infringing or unsafe). Retraining T2I models after filtering out unacceptable concepts in the training data is inefficient and degrades utility. Hence, we need concept removal techniques (CRTs) which are i) effectivein preventing the generation of images with unacceptable concepts ii) utility-preservingon acceptable concepts, and iii) robustagainst evasion with adversarial prompts. None of the prior CRTs satisfy all these requirements simultaneously.
@@
We introduce Espresso, the first robust concept filter based on Contrastive Language-Image Pre-Training (CLIP). We configure CLIP to identify unacceptable concepts in generated images using the distance of their embeddings to the text embeddings of both unacceptable and acceptable concepts. This simple change significantly improves robustness by restricting the adversary to add noise only along the vector connecting the embeddings of unacceptable and acceptable concepts. Further, fine-tuning Espresso to separate embeddings of unacceptable and acceptable concepts, while preserving their pairing with image embeddings, improves utility. We evaluate Espresso on eleven concepts to show that it is more effective (CLIP accuracy on unacceptable concepts in [0.00, 0.20] and FNR in [0.14, 0.53]) and more robust (CLIP accuracy on adversarial prompts for unacceptable concepts in [0.00, 0.40] and FNR in [0.00, 0.40]) than prior CRTs, while retaining utility (normalized CLIP score on acceptable concepts in [0.59, 0.98] and FPR in [0.01, 0.08]). Finally, we present theoretical bounds for certified robustness of Espresso against adversarial prompts, and an empirical analysis.
{arxiv}§§footnotetext: Work done while visiting the Secure Systems Group, University of Waterloo.
1 Introduction
@@
Diffusion based text-to-image (T2I) models have demonstrated a remarkable ability to generate high quality images from textual prompts [49, 50, 47]. These models are trained on large datasets of unfiltered content from the Internet [46, 53]. Due to their large capacity, T2I models memorize specific concepts, as seen in the generated images [28, 55, 6]. Some of these concepts, may be unacceptable for various reasons, such as copyright infringement (e.g., a movie character or celebrity), or inappropriateness (e.g., “nudity” or “violence”) [16, 52, 21]. Retraining T2I models after filtering out unacceptable concepts is inefficient, only partially effective [34, 51], and compromises utility [16]. Hence, there is a need for concept removal techniques (CRTs) to minimize unacceptable concepts in generated images.
@@
Ideally, CRTs should be effective in reducing the generation of unacceptable concepts while preserving the utility on all others, and robust to evasion with adversarial prompts. As we show in Section 6, none of the existing CRTs simultaneously satisfy these requirements: i) filtering CRTs, which detect unacceptable concepts, lack robustness ([48] and Section 6.2) ii) fine-tuning CRTswhich modify T2I models, trade-off effectiveness and utility [16, 30, 21, 52], and may lack robustness [56, 60, 42, 66]. Designing a CRT meeting all the requirements is an open problem. Our goal is to design such a CRT.
@@
We opt to use a filter as it will not alter the T2I model, thus reducing its impact on utility. We construct our filter using the Contrastive Language-Image Pre-Training (CLIP) [46], an essential component of T2I models. CLIP co-locates the embeddings of textual prompts and their corresponding images within a unified embedding space where similar concepts are closer. CLIP is pre-trained on a vast dataset which encodes a broad spectrum of concepts [46], making it a versatile choice for a filter, unlike specialized classifiers (e.g., [67]). However, relying on CLIP to identify unacceptable images by solely measuring the distance between the embeddings of a generated image, and an unacceptable concept, was shown not to be robust ( [48]). Subsequently, several prior works have identified filtering as a viable direction which needs further exploration [48, 34].
@@
We present a robust (CLIP-based) content filter, Espresso, by configuring CLIP to identify unacceptable concepts in generated images using the distance of their embeddings to the text embeddings of both unacceptable and acceptable concepts. This simple change makes it harder to generate effective adversarial prompts: the adversary is restricted to adding noise in the direction of the acceptable concept. To restore any drop in utility stemming from this change, we fine-tune Espresso to increase the separation between the text embeddings of unacceptable and acceptable concepts, while maintaining their pairing with their corresponding image embeddings. This fine-tuning approach is further supported by fine-tuning CRTs which show that such an optimization leads to better utility [30, 56, 15]. We claim the following main contributions: we present
  1. 1.
@@
    Espresso  [^1] We will open-source the code upon publication., the first robust (CLIP-based) content filter which identifies unacceptable concepts in generated images by measuring the distance of their embeddings to the text embeddings of both unacceptable and acceptable concepts (Section 4),
  2. 2.
@@
    a comprehensive comparative evaluation (Section 5) of the fine-tuned variant of Espresso with six state-of-the-art fine-tuning CRTs, and one filtering CRT, showing that it is more effective (CLIP accuracy on unacceptable concepts in [0.00, 0.20] and FNR in [0.14, 0.53]) and more robust (CLIP accuracy on adversarial prompts for unacceptable concepts in [0.00, 0.40] and FNR in [0.00, 0.40]) than prior CRTs, while retaining utility (normalized CLIP score on acceptable concepts in [0.59, 0.98] and FPR in [0.01, 0.08]). (Section 6), and
  3. 3.
@@
2 Background
We describe T2I models (Section 2.1), different CRTs (Section 2.2), and attacks against them (Section 2.3).
2.1 Diffusion based T2I Models
@@
where $\tau$ is the temperature parameter for scaling the predictions [46, 35].
@@
Several prominent T2I models (e.g., Stable Diffusion v1.4 (SDv1.4)) improve the efficiency of the diffusion process by conducting it in the embedding space of a variational autoencoder (VAE), rather than on images [49]. For a VAE decoder $\mathit{D}$, VAE encoder $\mathcal{E}$, and $z_{t}\in\mathcal{E}(\texttt{$x$})$ as the latent representation of $x$ in the VAE’s latent space, $\epsilon_{\theta}$ is trained by minimizing the following objective:
$\mathcal{L}=\mathbb{E}_{\epsilon,z_{t},\phi_{p}(\texttt{p}),t}[||\epsilon-\epsilon_{\theta}(z_{t},\phi_{p}(\texttt{p}),t)||_{2}^{2}]$ where $\epsilon\sim\mathcal{N}(0,1)$. The final image is generated from the approximation, $\tilde{z_{0}}$, by passing it through $\mathit{D}$: $\tilde{x_{0}}=\mathit{D}(\tilde{z_{0}})$. Table XI in Appendix .4 summarizes frequently used notations.
2.2 Concept Removal Techniques
The textual phrase for an acceptable concept is $c^{a}$, and for an unacceptable concept is $c^{u}$. For a given $c^{u}$, $c^{a}$ is either the opposite (e.g., $c^{u}$ = violence vs. $c^{a}$ = peaceful) or a general category of a specific character/object/person (e.g., $c^{u}$ = R2D2 vs. $c^{a}$ = robot). We discuss the selection of $c^{a}$ for a given $c^{u}$ in Section 5.3. An image $x$ generated from a T2I model may either contain an unacceptable concept (referred to as $x^{u}$) or an acceptable one (referred to as $x^{a}$). Similarly, a text prompt p may contain a phrase for an acceptable concept ($\texttt{p}^{a}$) or an unacceptable concept ($\texttt{p}^{u}$). An example of an unacceptable prompt $\texttt{p}^{u}$ containing an unacceptable concept $c^{u}$ = Captain Marvel, is “Captain Marvel soaring through the sky”. CRTs seek to thwart the generation of $x^{u}$ by either fine-tuning the T2I model to suppress $x^{u}$, or using a classifier as a filter to detect $x^{u}$, and serve a replacement image instead. We first present six state-of-the-art fine-tuning CRTs:
Concept Ablation (CA[30] fine-tunes the T2I model to minimize the KL divergence between the model’s output for $\texttt{p}^{u}$ and $\texttt{p}^{a}$ to force the generation of $x^{a}$ instead of $x^{u}$. Formally, they optimize the following objective function:
$\begin{split}\footnotesize\mathcal{L}_{\texttt{CA}}=\mathbb{E}_{\epsilon,z_{t},c^{u},c^{a},t}[w_{t}||\epsilon_{\theta}(z_{t},\phi_{p}(\texttt{p}^{a}),t).sg()\\ -\epsilon_{\theta}(z_{t},\phi_{p}(\texttt{p}^{u}),t)||_{2}^{2}]\end{split}$
@@ 2.2 Concept Removal Techniques
Forget-Me-Not (FMN[68] minimizes the activation maps for $c^{u}$ by modifying $\epsilon_{\theta}$’s cross-attention layers. Further, fine-tuning $\epsilon_{\theta}$, instead of just the cross-attention layers, results in degraded utility.
Selective Amnesia (SA[21] fine-tunes T2I models by adapting continuous learning techniques (elastic weight consolidation and generative replay) for T2I models to forget a concept. They optimize $\mathbb{P}(\texttt{$x$}|\theta^{*},c^{u})$, the probability of generating $x$ given $\theta^{*}$ and $c^{u}$, where $\theta^{*}$ are the frozen parameters of the original T2I model:
@@
$\begin{split}\footnotesize\mathcal{L}_{\texttt{SA}}=-\mathbb{E}_{\mathbb{P}(\texttt{$x$}|\texttt{p})\mathbb{P}_{f}(c^{u})}[\text{log}\mathbb{P}(\texttt{$x$}|\theta^{*},c^{u})]-\lambda\sum_{i}\frac{M_{i}}{2}(\theta_{i}^{*}-\theta_{i})^{2}\\ +\mathbb{E}_{\mathbb{P}(\texttt{$x$}|\texttt{p})\mathbb{P}_{r}(c^{a})}[\text{log}\mathbb{P}(\texttt{$x$}|\theta,c^{a})]\end{split}$
Erased Stable Diffusion (ESD[16] fine-tunes the T2I model by modifying the reverse diffusion process to reduce the probability of generating $x^{u}$:
$\begin{split}\epsilon_{\theta}(\texttt{$x$}_{t},\phi_{p}(c^{u}),t)=\epsilon_{\theta^{*}}(\texttt{$x$}_{t},t)+\eta(\epsilon_{\theta^{*}}(\texttt{$x$}_{t},\phi_{p}(c^{u}),t)\\ -\epsilon_{\theta^{*}}(\texttt{$x$}_{t},t))\end{split}$
Unified Concept Editing (UCE[19] fine-tunes the T2I model’s cross-attention layers with a language model to minimize the influence of $c^{u}$, while keeping the influence of remaining concepts unchanged. Their optimization is:
$\begin{split}\mathcal{L}_{\texttt{UCE}}=\sum_{c^{u}\in\mathcal{C}^{u},c^{a}\in\mathcal{C}^{a}}||W\times c^{u}-W^{*}\times c^{a}||_{2}^{2}\\ +\sum_{c\in S}||W\times c-W^{*}\times c||_{2}^{2}\end{split}$
where $W$, $W^{*}$ are the parameters of the fine-tuned and original cross-attention layers in $\epsilon_{\theta}$, $\mathcal{C}^{u}$ and $\mathcal{C}^{a}$ are the space of pre-defined unacceptable and acceptable concepts, and $S$ is a set of concepts for which to preserve utility.
Safe diffusion (SDD[26] fine-tunes the T2I model by encouraging the diffusion model noise conditioned on $c^{u}$ to match the unconditioned noise, while minimizing the utility drop using the following objective function:
$\begin{split}\mathcal{L}_{\texttt{SDD}}=\mathbb{E}_{\epsilon,z_{t},c^{u},t}[||\epsilon_{\theta}(z_{t},\phi_{p}(c^{u}),t)-\epsilon_{\theta}(z_{t},t).sg()||_{2}^{2}]\end{split}$
where $\epsilon\sim\mathcal{N}(0,1)$ and $z_{t}\in\mathcal{E}(x)$.
We now describe two filtering CRTs:
Stable Diffusion Filter (SD-Filter) [46] is black-box and the design of the filter is not publicly available. However, Rando et. al. [48] hypothesize that it involves computing the cosine similarity between the embeddings of a generated image, $x$, and a pre-defined set of $c^{u}$. If the cosine similarity is greater than some threshold ($\Gamma$), then $x$ has $c^{u}$. Formally, their filter $\texttt{F}_{SD}$ can be described as
where $\text{cos}(\phi_{x}(\texttt{$x$}),\phi_{p}(c^{u}))=\overline{\phi_{x}(\texttt{$x$})}\cdot\overline{\phi_{p}(c^{u})}$, and $\overline{\cdot}$ denotes normalization. Here $\texttt{F}_{SD}(\texttt{$x$})=0$ indicates $x^{a}$ and $\texttt{F}_{SD}(\texttt{$x$})=1$ indicates $x^{u}$. Note that $\Gamma$ varies with $c^{u}$.
Unsafe Diffusion (UD[67] is the current state-of-the-art filtering CRT and outperforms SD-Filter [67]. UD trains a multi-headed neural network classifier on top of CLIP to identify $x^{u}$ where each head classifies different $c^{u}$: nudity, violence, disturbing, hateful, and political. Their objective is given as $\texttt{F}_{\texttt{UD}}(\texttt{$x$})=\text{MLP}(\phi_{x}(\texttt{$x$}))$ where MLP is a multi-layer perceptron, and $\texttt{F}_{\texttt{UD}}\in\{0,1\}$.
2.3 Evading Concept Removal Techniques
@@
PEZ [60] constructs $\texttt{p}^{adv}$ by identifying text tokens by minimizing: $\mathcal{L}_{PEZ}=1-\text{cos}(\phi_{p}(\texttt{p}^{adv}),\phi_{x}(\texttt{$x^{u}$}))$.
RingBell [56] identifies tokens for $\texttt{p}^{adv}$ by first calculating the average difference between the embedding vectors of $\texttt{p}^{u}$ and $\texttt{p}^{a}$: $\phi_{p}(\hat{\texttt{p}})=\frac{1}{N}\sum_{i=1}^{N}\{\phi_{p}(\texttt{p}^{u}_{i})-\phi_{p}(\texttt{p}^{a}_{i})\}$ from a set of $N$ prompts. Then, RingBell computes $\texttt{p}^{adv}$ by minimizing the distance of its text embedding to ($\phi_{p}(\texttt{p}^{init})+\eta\cdot\phi_{p}(\hat{\texttt{p}}))$, formulated as: $\text{min}_{\texttt{p}^{adv}}||\phi_{p}(\texttt{p}^{adv})-(\phi_{p}(\texttt{p}^{init})+\eta\cdot\phi_{p}(\hat{\texttt{p}}))||^{2}$ where $\texttt{p}^{init}$ is an initial unacceptable prompt which is censored by a CRT. They solve this using a genetic algorithm.
SneakyPrompt [66] uses reinforcement learning to construct $\texttt{p}^{adv}$ specifically against filtering CRTs. Given an initial prompt $\texttt{p}^{init}$, the attack searches for tokens to update $\texttt{p}^{init}$ to form $\texttt{p}^{adv}$. The reward function is the cosine similarity between $\phi_{x}(\texttt{$x$}^{adv})$ and $\phi_{p}(\texttt{p}^{adv})$. Typically, they use $\texttt{p}^{u}$ as $\texttt{p}^{init}$ for better effectiveness and faster convergence.
CCE [42] uses textual-inversion to construct $\texttt{p}^{adv}$ [14]. $\mathcal{A}dv$ updates CLIP’s vocabulary to include a new token “<s>” which, when included in $\texttt{p}^{adv}$, generates $\mathcal{A}dv$’s desired image. To optimize <s>, we find $v$, an embedding for <s>, corresponding to least loss $\mathcal{L}_{CCE}=\mathcal{L}_{MSE}(\epsilon,\epsilon_{\theta}(z_{t},v,t))$, where $\epsilon\sim\mathcal{N}(0,1)$, and $\mathcal{L}_{MSE}$ is the mean-squared-error.